OpenStudy (anonymous):

I need help.. First estimate the answer. Then divide to find the exact answer. Simplify all answers. 19/10 divide by 4 3/5 = The estimate answer is_____ The exact answer is______

OpenStudy (cwrw238):

19/10 divide by 4 3/5 = approximately 2 / 5 4 3/5 = 23 / 5 19/10 / 23/5 = 19/10 * 5 / 23

OpenStudy (anonymous):

im confuse here. so what is the estimate answer

OpenStudy (cwrw238):